Could you help me with 11,12,and 13 please ?

Thanks in advance,

Massimo Ulto8. If you have had calculus, prove the power rule for positive exponents. Specifically, prove that for every positive integer
Answer #1


A-11 for n= 1 2 +2 3+---+n? = (n (n+1)/ ) use Induction Method - *** 13 = 44.() = 1.1 Now , assume the formula is true for n= 13 for every positive integer 2 > n Induction method for n= 2 1 1 > 1 which true. is Assume it is true for nam je, am>m We w